# Nonlinearity-induced Thouless pumping of solitons

### v1

@@ Abstract
It has recently been theoretically predicted and experimentally observed that a soliton resulting from nonlinearity can be pumped across an integer or fractional number of unit cells as a system parameter is slowly varied over a pump period. Nonlinear Thouless pumping is now understood as the flow of instantaneous Wannier functions, ruling out the possibility of pumping a soliton across a nonzero number of unit cells over one cycle when a corresponding Wannier function does not exhibit any flow, i.e., when the corresponding Bloch band that the soliton bifurcates from is topologically trivial. Here we surprisingly find an anomalous nonlinear Thouless pump where the displacement of a soliton over one cycle is twice the Chern number of the Bloch band from which the soliton bifurcates. We show that the breakdown of the correspondence between the displacement of a soliton and the Chern number arises from the emergence of a new branch of stable soliton solutions mainly consisting of two neighboring instantaneous Wannier functions. Furthermore, we find a nonlinearity-induced integer quantized Thouless pump, allowing a soliton to travel across one unit cell during a pump period, even when the corresponding band is topologically trivial. Our results open the door to studying nonlinearity-induced Thouless pumping of solitons.
@@
Thouless pumping, which plays an important role in understanding the quantum Hall effect, is a phenomenon where quantized transport arises through the slow periodic variation of a Hamiltonian parameter [1]. The transport requires the complete filling of an energy band so that it is dictated by the Chern number of the occupied band with the varying parameter playing the role of an additional momentum. So far, Thouless pumps have been experimentally observed in photonic systems [2, 3, 4], cold atom systems [5, 6, 7, 8, 9] and other systems [10, 11, 12].
@@
Nonlinearity is a natural feature in various systems, including photonic systems [13, 14, 15, 16, 17, 18, 19] and Bose-Einstein condensates (BECs) [20, 21, 22, 23, 24, 25]. It leads to many intriguing phenomena, such as solitons, which are wave packets that travel without changing their shape. Nonlinearity also gives rise to topology-related phenomena, such as topological bulk [26, 27] and edge solitons [28, 29, 30, 31], as well as nonlinearity-induced topological insulators [32, 33]. In particular, it has recently been found that a soliton arising from nonlinearity can exhibit quantized transport when an underlying Hamiltonian parameter is slowly tuned periodically (called nonlinear Thouless pumping) [34, 35, 36, 37, 38, 39]. Later, a quantized fractional Thouless pumping of solitons is found by considering a soliton bifurcating from multiple bands [40]. The integer and fractional quantized Thouless pumping are now understood as the flow of the instantaneous one-band and maximally localized multi-band Wannier functions [35, 36, 37, 40], respectively. For the former, the displacement of a soliton is determined by the Chern number of the Bloch band of the linear Hamiltonian of the system [35, 36, 37]. If this mechanism always worked, we would not observe the occurrence that a soliton’s displacement is larger than this Chern number. In other words, a soliton would not undergo transport when a linear Hamiltonian is topologically trivial.
@@
Here we surprisingly find an anomalous nonlinear Thouless pump that exhibits the pumping of a soliton across two unit cells over one cycle, while the Chern number of the band, from which the soliton bifurcates, is $C=-1$. Our result is thus beyond the previous nonlinear Thouless pump where the displacement is identical to the Chern number. It cannot be attributed to the renormalization of a system parameter resulting from nonlinearity, as variations of parameters only permit the Chern number to change from $-1$ to $1$, rather than to $2$ in our system. We show that the phenomenon arises due to the emergence of a new branch of stable instantaneous soliton solutions formed by a superposition of two neighboring instantaneous Wannier functions. These soliton solutions enable the displacement to be two times the Chern number. Leveraging the discovery of the new nonlinear Thouless pump, we construct an alternative nonlinear model, demonstrating the occurrence of quantized Thouless pumping even when the Chern number of the band of the linear Hamiltonian vanishes (in other words, the linear Thouless pumping does not happen in this system over one cycle). To further characterize these anomalous phenomena, we study a spatial supercell containing multiple unit cells to ensure the well localization of a soliton within the supercell. Subsequently, we calculate the Chern number of a modulated Hamiltonian involving effects of the soliton solution, finding it to be identical to the displacement. Finally, we show the emergence of anomalous nonlinear Thouless pumping in a continuous model, which can be experimentally realized in cold atom systems.
Nonlinear model.— To illustrate our findings, we start by investigating the following dimensionless discrete nonlinear Schrödinger equation:
$i\frac{\partial}{\partial t}\psi_{\sigma x}=$ $\sum_{\sigma^{\prime}x^{\prime}}H_{\sigma x,\sigma^{\prime}x^{\prime}}^{\textrm{lin}}(\theta)\psi_{\sigma^{\prime}x^{\prime}}$ (1)
$+\left(g|\psi_{\sigma x}|^{2}+g_{12}|\psi_{\overline{\sigma}x}|^{2}\right)\psi_{\sigma x},$
where $\psi_{\sigma x}$ is the wavefunction of the $\sigma$th component ($\sigma=1,2$) in the $x$th unit cell at time $t$ ($t$ is the propagation distance in photonic systems), $H^{\textrm{lin}}$ is the linear tight-binding Hamiltonian that depends on a real system parameter $\theta$, $g$ and $g_{12}$ are nonlinear coefficients describing the intraspecies and interspecies interactions in ultracold atomic gases, respectively [$\overline{\sigma}=(\sigma\text{ mod }2)+1$ labels the other component]. The norm of the wavefunction $N=\sum_{\sigma x}|\psi_{\sigma x}|^{2}$ is preserved during the evolution in this equation. To show that the displacement of a soliton can be twice the Chern number of a Bloch band, we choose a linear Hamiltonian which reads in momentum space
@@
where $\sigma_{\nu}$ ($\nu=x,y,z$) are Pauli matrices, and $m_{z}=m_{0}+\cos\theta$ with $m_{0}$ being a real system parameter. We set $J_{1}=J_{1}^{\prime}=1$ and $J_{2}=\sin\theta$. The parameter $\theta$ is slowly varied as time evolves, realizing a periodic change of the linear Hamiltonian. Clearly, this model realizes the Chern band [41] if we view $\theta$ as the other momentum $k_{y}$ besides $k$. The Chern number of the lowest band with respect to $k$ and $\theta$ is $C=1$ when $-2<m_{0}<0$, $C=-1$ when $20<m_{0}<$ 2, and $C=0$ otherwise. In the following study concerning a soliton’s displacement of $-2$, we take $m_{0}=1$. For linear Thouless pumping, particles are required to fill the entire band so that the cloud of particles will travel across the lattice distance identical to the Chern number as we slowly tune $\theta$ from 00 to $2\pi$ [1]. Note that this tight-binding model can be experimentally implemented in ultracold atomic gases as proposed in Ref. [42] (see the following discussion).
Figure 1: Anomalous nonlinear Thouless pumping. a, One-cycle evolution of the density distribution $|\psi_{x}|^{2}=\sum_{\sigma}|\psi_{\sigma x}|^{2}$ of instantaneous solitons which bifurcate from the lowest band when $g_{12}=0$. b, The trajectory of center-of-mass positions of instantaneous solitons when $g_{12}=0$ (red line), $g_{12}=1$ (blue line) and the corresponding Wannier functions (black line) as a system parameter $\theta$ varies. Here, we set $N=1.45$ and $g=1$.
Nonlinear evolution and mechanism.— With nonlinearity, spatially localized solutions known as solitons can emerge due to the balance between nonlinearity and dispersion [17, 19, 22, 23]. Remarkably, previous studies have demonstrated that in a nonlinear Aubry-André-Harper model, a soliton can be transported across an integer number of unit cells per cycle, which is dictated by the Chern number of the underlying Bloch band from which the soliton bifurcates [34]. The phenomenon is now understood as the motion of a soliton following the center-of-mass positions (referred to as the Wannier center) of Wannier functions of the corresponding Bloch bands [34, 35, 36, 37, 38, 40]. In our case, given that the Chern number of the lowest band is $C=-1$, we expect that the displacement of a soliton per cycle should be $-1$ (in units of the length of a unit cell, which we will take hereafter) based on the theory. However, contrary to our expectations, we find that the actual displacement is $-2$, deviating from the value predicted by the Chern number.
@@
To demonstrate our findings, we calculate stable instantaneous soliton solutions that bifurcate from the lowest band with the Chern number $C=-1$ for the instantaneous nonlinear Hamiltonian at each $\theta$ using the Newton’s method. This approach is justified because, in the adiabatic limit—where $\theta$ is varied very slowly—the time evolution of a soliton follows the instantaneous eigenstate of the instantaneous nonlinear Hamiltonian at time $t$ [43, 44, 45, 46, 47]. We have directly computed the time evolution based on Eq. (1) when $\theta$ is varied sufficiently slowly and found that the results closely resemble the instantaneous solutions. Additionally, we have performed a stability analysis to confirm that the solved nonlinear instantaneous soliton solutions remain stable at each $\theta$ (see also Supplementary Information Section S-1).
Figure 1b illustrates that when $g_{12}=1$, the soliton travels across the lattices by a distance of $-1$ over one period, which is identical to the Chern number of the lowest band. The result is consistent with the previous findings [34, 35, 36, 37]. Suprisingly, we observe that when $g_{12}=0$, the soliton’s displacement becomes $-2$, twice the Chern number of the lowest band, as shown in Fig. 1. We find that the transition between the two displacements occurs near $g_{12}=0.2$ when $N=1.45$. The phenomenon cannot be attributed to the renormalization of the system parameter $m_{0}$ resulting from nonlinearity, as variations in $m_{0}$ only enable the Chern number to change from $-1$ to $1$, rather than to $2$.
@@
Figure 2: Mechanism of anomalous nonlinear pumping. a-d, Center-of-mass positions of stable instantaneous solitons (black and red lines) and Wannier functions (dashed grey lines) with respect to a system parameter $\theta$ over one cycle for different $g_{12}$. The anomalous nonlinear pumping arises due to the appearance of a new branch of soliton solutions around $\theta=\pi$ plotted as red lines in b. Here, we set $N=1.45$ and $g=1$.
@@
To explain the phenomenon, we plot in Fig. 2 the center-of-mass position of stable nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as a function of $\theta$ for different values of $g_{12}$. Notably, when $g_{12}=0.3$, the center-of-mass positions of a soliton and the Wannier centers exhibit apparent differences for most values of $\theta$, but they coincide at $\theta=0$ and $\theta=\pi$. This results in a displacement of $-1$ over each period, following the trajectory indicated by the black lines. In fact, at $\theta=\pi$, the soliton solution is exactly identical to the Wannier function multiplied by $\sqrt{N}$ when $m_{0}=1$, as will be further elucidated below. However, as $g_{12}$ is decreased, a stable soliton solution emerges at $\theta=\pi$, localized around $x=n$ with $n$ being an integer, differing from the Wannier center, which is located at $n+1/2$. With a further decrease in $g_{12}$, we observe that the new branch of solutions becomes longer, eventually connecting to the original soliton solutions (represented by black lines). Consequently, when $g_{12}<0.2$, a soliton becomes displaced by $-2$ unit cells during adiabatic evolution, following the trajectory indicated by the black lines in Fig. 2c,d.
To elucidate the emergence of the nonlinear solution at $\theta=\pi$ localized around $x=n$, we transform the linear Hamiltonian to the Su-Schrieffer-Heeger (SSH) model, expressed as, $H^{\textrm{lin}}_{\textrm{SSH}}=UH^{\textrm{lin}}U^{\dagger}=(m_{z}+\cos k)\sigma_{x}+\sin k\sigma_{y}$ with $U=e^{-i\pi\sigma_{y}/4}$. When $m_{0}=1$ so that $m_{z}=0$, the model in real space contains only intercell hopping. In this scenario, we show that the SSH model’s Wannier function multiplied by $\sqrt{N}$ is a solution to the transformed nonlinear equation (see Supplementary Information Section S-2). However, we find that these solutions become unstable when $g_{12}$ is small.
For solutions at $\theta=\pi$ that localize between two adjacent Wannier functions, it is likely that these solutions represent a superposition of these Wannier functions. Indeed, we find that these solutions mainly consist of equal superpositions of two neighboring Wannier functions of the lowest band (see Supplementary Information Section S-2). We note that similar solutions have also been found in Ref. [35], but these solutions are unstable in their case. In contrast, although our solutions are unstable for large $g_{12}$, they become stable for smaller values of $g_{12}$, facilitating the anomalous transport of a soliton. While we focus on the specific case of $m_{0}=1$ for simplicity of analysis, we have confirmed that anomalous nonlinear Thouless pumping occurs for other values of $m_{0}$ as well.
@@
Nonlinearity-induced Thouless pumping.— We have demonstrated the breakdown of the correspondence between displacement and the Chern number for a nonlinear Thouless pump, which naturally leads to a question whether a soliton can be pumped when the Chern number of the linear band is zero. Our answer is affirmative. To illustrate this, we again consider the linear Hamiltonian in Eq. (2). Now, we vary the system parameters $m_{z}$, $J_{1}=J_{1}^{\prime}$ and $J_{2}$ with respect to $\theta$, as depicted in Fig. 3(a). For simplicity, we impose the conditions that $m_{z}(\theta)=m_{z}(2\pi-\theta)$, $J_{1}(\theta)=J_{1}(2\pi-\theta)$ and $J_{2}(\theta)=J_{2}(2\pi-\theta)$, ensuring that these parameters are symmetric about $\theta=\pi$. Consequently, the Hamiltonian $H^{\textrm{lin}}(k)$ also exhibits symmetry with respect to $\theta$, i.e., $H^{\textrm{lin}}(k,\theta)=H^{\textrm{lin}}(k,2\pi-\theta)$. The immediate result is that the Berry phase of the $n$th ($n=1,2$) band of $H^{\textrm{lin}}(k,\theta)$ satisfies $\gamma_{n}(\theta)=\gamma_{n}(2\pi-\theta)$ [where $\gamma_{n}(\theta)$ denotes the Berry phase following the path along $k$ for the $n$th band at a fixed $\theta$], leading to a Chern number of zero for each band. Therefore, the linear Thouless pumping cannot occur.
@@
To achieve nonlinearity-induced Thouless pumping of solitons, we simultaneously adjust $g_{12}$, as shown in Fig. 3b (the interaction can be controlled via Feshbach resonances in ultracold atomic gases [24]). Specifically, we initially reduce $g_{12}$ to zero gradually, which facilitates the soliton’s travel over a longer distance. Subsequently, we slowly increase $g_{12}$ back to its original value to prevent the soliton from returning to its initial position. In Fig. 3c, we plot the density distribution of nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as $\theta$ varies from 00 to $2\pi$, remarkably demonstrating that the soliton is displaced by $-1$ unit cell. This effect is further evidenced in Fig. 3d, which depicts the evolution of center-of-mass positions of instantaneous soliton solutions as $\theta$ changes. This observation starkly contrasts with the behavior of the Wannier center (or the Berry phase), which returns to its starting value after one complete cycle. Hence, we successfully achieve Thouless pumping of a soliton through the influence of nonlinearity.
@@
Figure 3: Nonlinearity-induced Thouless pumping. a,b, Illustration of how the parameters $m_{z}(\theta)$, $J_{1}(\theta)$, $J_{2}(\theta)$ and $g_{12}(\theta)$ are varied with respect to $\theta$ in order to induce a nonlinear pumping for a topologically trivial band. c, One-cycle evolution of the density distribution $|\psi_{x}|^{2}$ of instantaneous solitons bifurcating from the lowest band. d, The evolution of center-of-mass positions of instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one cycle. Here, we set $N=1.25$ and $g=1$.

### v2

# Anomalous quantized nonlinear Thouless pumping

@@ Abstract
It has recently been theoretically predicted and experimentally observed that a soliton resulting from nonlinearity can be pumped across an integer or fractional number of unit cells as a system parameter is slowly varied over a pump period. Nonlinear Thouless pumping is now understood as the flow of instantaneous Wannier functions, ruling out the possibility of pumping a soliton across a nonzero number of unit cells over one cycle when a corresponding Wannier function does not exhibit any flow, i.e., when the corresponding Bloch band that the soliton bifurcates from is topologically trivial. Here we surprisingly find an anomalous nonlinear Thouless pump where the displacement of a soliton over one cycle differs from the Chern number of the Bloch band from which the soliton comes. We develop a general theory showing that this anomalous behavior arises from a transition of a soliton between different Wannier functions by passing through an intersite-soliton (or dipole-soliton) state. Furthermore, we find a nonlinearity-induced integer quantized Thouless pump of a soliton, allowing a soliton to travel across one unit cell during a pump period, even when the corresponding band is topologically trivial. Our results open the door to studying nonlinearity-induced Thouless pumping of solitons.
@@
Thouless pumping, which plays an important role in understanding the quantum Hall effect, is a phenomenon where quantized transport arises through the slow periodic variation of a Hamiltonian parameter [1, 2]. The transport requires the complete filling of an energy band so that it is dictated by the Chern number of the occupied band with the varying parameter playing the role of an additional momentum. So far, Thouless pumps have been experimentally observed in photonic systems [3, 4, 5], cold atom systems [6, 7, 8, 9, 10] and other systems [11, 12, 13].
@@
Nonlinearity is a natural feature in various systems, including photonic systems [14, 15, 16, 17, 18, 19, 20] and Bose-Einstein condensates (BECs) [21, 22, 23, 24, 25, 26]. It leads to many intriguing phenomena, such as solitons, which are wave packets that travel without changing their shape. Nonlinearity also gives rise to topology-related phenomena, such as topological bulk [27, 28] and edge solitons [29, 30, 31, 32], as well as nonlinearity-induced topological insulators [33, 34]. In particular, it has recently been found that a soliton arising from nonlinearity can exhibit quantized transport when an underlying Hamiltonian parameter is slowly tuned periodically (called nonlinear Thouless pumping)  [35, 36, 37, 38, 39, 40, 41]. Later, a quantized fractional Thouless pumping of solitons is found by considering a soliton bifurcating from multiple bands [42]. The integer and fractional quantized Thouless pumping are now understood as the flow of the instantaneous one-band and maximally localized multi-band Wannier functions [36, 37, 38, 42], respectively. For the former, the displacement of a soliton is determined by the Chern number of the Bloch band of the linear Hamiltonian of the system [36, 37, 38]. If this mechanism always worked, we would not observe the occurrence that a soliton’s displacement is different from this Chern number. In other words, a soliton would not undergo transport when a linear Hamiltonian is topologically trivial.
@@
Here we surprisingly find an anomalous nonlinear Thouless pump that can exhibit the pumping of a soliton across zero, two, or three unit cells over one cycle, while the Chern number of the band, from which the soliton comes, is C=1C=-1. Our result is thus beyond the previous nonlinear Thouless pump where the displacement is identical to the Chern number. We develop a general theory revealing that this anomalous behavior arises from a transition of a soliton between different Wannier functions by passing through an intersite-soliton (or dipole-soliton) state. Leveraging the discovery of the new nonlinear Thouless pump, we construct an alternative nonlinear model, demonstrating the occurrence of quantized Thouless pumping even when the Chern number of the band of the linear Hamiltonian vanishes (in other words, the linear Thouless pumping does not happen in this system over one cycle). To further characterize the topological properties of these anomalous phenomena, we study a spatial supercell containing multiple unit cells to ensure the well localization of a soliton within the supercell. Subsequently, we calculate the Chern number of a modulated Hamiltonian involving effects of the soliton solution, finding it to be identical to the displacements. Finally, we show the emergence of anomalous nonlinear Thouless pumping in a nonlinear continuous model, which can be experimentally realized in cold atom systems.
General theory.—We start by proposing a general theory showing how the anomalous nonlinear Thouless pump can occur. Consider a one-dimensional (1D) non-interacting Hamiltonian $H^{\mathrm{lin}}(\theta)$ with translation symmetry. The Hamiltonian is controlled by a system parameter $\theta$ which varies from 0 to $2\pi$, satisfying $H^{\mathrm{lin}}(0)=H^{\mathrm{lin}}(2\pi)$. The Hamiltonian at each $\theta$ has Bloch states characterized by momentum $k$. We focus on an isolated Bloch band associated with a set of Wannier functions denoted by $W_{l}(j,\theta)$, where $l$ is an integer denoting the lattice vector in 1D, and $j$ is the spatial coordinate. Without loss of generality, we assume that the Chern number of this band with respect to momentum $k$ and parameter $\theta$ is 1-1 (although our theory is applicable to other Chern numbers as well). Consequently, the center-of-mass position (referred to as the Wannier center) of $W_{l}(j,\theta)$ changes by 1-1 as $\theta$ varies from 0 to $2\pi$, assuming that the length of a unit cell is $1$, as shown by the black lines in Fig. 1. Prior research indicates that, in the presence of nonlinearity, soliton solutions that resemble Wannier functions (referred to as onsite solitons) exist, such that the soliton undergoes a displacement equal to the Chern number over one pump period [35, 36, 37, 38]. Therefore, in this case, the soliton experiences a displacement of 1-1 over one cycle.
Figure 1: Schematic illustration of how anomalous nonlinear Thouless pumping arises. The black lines represent the center-of-mass position of Wannier functions with respect to $\theta$. When a soliton follows a black line, normal nonlinear pumping appears. In contrast, if a soliton transitions between Wannier functions by passing through intersite solitons, whose center-of-mass positions are marked by diagonal crosses, anomalous nonlinear pumping occurs. For example, following the blue and red lines in a results in displacements of 0 and 2-2, respectively, while following the green line in b results in a displacement of 3-3.
In addition to the onsite-soliton solution, previous studies have also investigated intersite (or dipole) soliton solutions [18, 19, 20]. We define an ideal intersite soliton as an equal superposition of two neighboring Wannier functions, e.g., $W_{0}(j,\theta)+W_{1}(j,\theta)]/\sqrt{2}$. Note that while the intersite soliton might be reminiscent of the doublon [43], they are completely different. Although intersite solitons themselves have been extensively studied [44, 45, 46, 47, 48, 49, 50, 51, 52], in the context of nonlinear Thouless pumping, prior studies have found that an intersite soliton is always unstable [36], thus precluding the stable evolution between onsite and intersite solitons, as a system parameter is varied.
We now show that a soliton can transition from one Wannier function to another by passing through an intersite-soliton state, leading to the anomalous nonlinear Thouless pumping. Specifically, consider a system where, at $\theta=0$, a soliton solution is $W_{0}(j,0)$, with a center-of-mass position of 0. If the soliton remains in the Wannier state as $\theta$ increases, its center-of-mass will follow the black trajectory, as illustrated in Fig. 1, consistent with previous reports. However, we find that near $\theta=\pi$, stable intersite solitons exist, e.g., $\left[W_{0}(j,\pi)+W_{1}(j,\pi)\right]/\sqrt{2}$ and $\left[W_{0}(j,\pi)+W_{-1}(j,\pi)\right]/\sqrt{2}$, with their center-of-mass positions marked by diagonal crosses in Fig. 1a. Through such an intersite soliton, the soliton can transition from $W_{0}(j,0)$ into either $W_{1}(j,2\pi)$ or $W_{-1}(j,2\pi)$, following the blue or red trajectory and yielding quantized transports of 0 and 2-2, respectively, both of which are distinct from the Chern number. Although the former case is similar to the trapped soliton for strong nonlinearity [35, 37], the mechanism is completely different (see Supplementary Information Section S-1 for details). The trapped phenomenon arises from all band contributions due to strong nonlinearity, leading to zero Chern number [42]. In contrast, in our case, the soliton always occupies one single band.
Moreover, as shown in Fig. 1b, a soliton can first evolve from $W_{0}(j,0)$ to $W_{-1}(j,\pi)$ through an intersite-soliton state formed by $W_{0}$ and $W_{-1}$ and subsequently evolve to $W_{-2}(j,2\pi)$ through an intersite-soliton state formed by $W_{-1}$ and $W_{-2}$. This process results in a displacement of 3-3. These anomalous transport outcomes do not align with the Chern number of the linear band. The theory suggests the potential for various anomalous results; however, it is crucial that the soliton solutions remain stable throughout the entire period.
Figure 2: Anomalous nonlinear Thouless pumping in the discrete nonlinear model. a, One-cycle trajectory of center-of-mass positions of instantaneous solitons coming from the lowest band of the linear Hamiltonian in Eq. (1) as a system parameter $\theta$ varies. It plots three anomalous cases corresponding to displacements of 0 (case 1), 2-2 (case 2), and 3-3 (case 3), described by the blue (g=1g=-1, $g_{12}=0$, and $m_{0}=1$), red ($g=1$, $g_{12}=0$, and $m_{0}=1$), and green lines ($g=1$, $g_{12}=0$, and $m_{0}=1.3$), respectively, in contrast to the normal case with a displacement of 1-1 described by the black line ($g=g_{12}=m_{0}=1$). The associated density distributions $|\psi_{j}|^{2}=\sum_{\sigma}|\psi_{\sigma j}|^{2}$ of the solitons for the three anomalous cases are plotted in b-d. e-h, Occupations $\rho_{W_{l}}(\theta)$ of instantaneous solitons on Wannier functions $W_{l}(\theta)$ of the lowest band along the Wannier centers as a function of $\theta$ for the four cases. See also the Supplementary Information Section S-3 for more information. Here, we set $N=1.45$.
Discrete nonlinear model.—To demonstrate that the anomalous nonlinear Thouless pumping can arise in concrete models, we study the following dimensionless discrete nonlinear Schrödinger equation:
$i\frac{\partial}{\partial t}\psi_{\sigma j}=\sum_{\sigma^{\prime}j^{\prime}}H_{\sigma j,\sigma^{\prime}j^{\prime}}^{\textrm{lin}}(\theta)\psi_{\sigma^{\prime}j^{\prime}}+{V_{\sigma j}(\psi_{\sigma j},\psi_{\overline{\sigma}j})\psi_{\sigma j}},$ (1)
[…]
@@
where $\sigma_{\nu}$ ($\nu=x,y,z$) are Pauli matrices, and $m_{z}=m_{0}+\cos\theta$ with $m_{0}$ being a real system parameter. We set $J_{1}=J_{1}^{\prime}=1$ and $J_{2}=\sin\theta$. The parameter $\theta$ is slowly varied as time evolves, realizing a periodic change of the linear Hamiltonian. This model can be transformed to the Rice-Mele model through a unitary transformation realized by $U=e^{-i\pi\sigma_{y}/4}$, and it clearly realizes the Chern band [53] if we view $\theta$ as the other momentum $k_{y}$ besides $k$. The Chern number of the lowest band with respect to $k$ and $\theta$ is $C=1$ when $2<m_{0}<0$, C=1C=-1 when $0<m_{0}<2$, and $C=0$ otherwise. For linear Thouless pumping, particles are required to fill an entire band so that the cloud of particles will travel across the lattice distance identical to the Chern number as we slowly tune $\theta$ from 0 to $2\pi$ [1]. Note that this tight-binding model can be experimentally implemented in ultracold atomic gases as proposed in Ref. [54] with $g$ and $g_{12}$ describing the intraspecies and interspecies interactions, respectively (see the section on nonlinear continuous model).
@@
To demonstrate the anomalous transport behavior in this nonlinear model, we calculate stable instantaneous soliton solutions $\chi_{\sigma j}(\theta)$ that come from the lowest band with the Chern number C=1C=-1 for the instantaneous nonlinear Hamiltonian at each $\theta$ using the Newton’s method; the solution is related to $\psi_{\sigma j}$ through $\psi_{\sigma j}=e^{-i\mu(\theta)t}\chi_{\sigma j}(\theta)$, where $\mu(\theta)$ is the chemical potential. This approach is justified because, in the adiabatic limit—where $\theta$ is varied very slowly—the time evolution of a soliton follows the instantaneous eigenstate of the instantaneous nonlinear Hamiltonian at time $t$ [55, 56, 57, 58, 59]. We have directly computed the time evolution based on Eq. (1) when $\theta$ is varied sufficiently slowly and found that the results closely resemble the instantaneous solutions. Additionally, we have performed a stability analysis to confirm that the solved nonlinear instantaneous soliton solutions remain stable at each $\theta$ (see also Supplementary Information Section S-2).
@@
Figure 3: Transition between normal and anomalous nonlinear Thouless pumping. a-d, Center-of-mass positions of stable instantaneous solitons (black and red lines) and Wannier functions (dashed grey lines) with respect to a system parameter $\theta$ over one cycle for different $g_{12}$. The anomalous nonlinear pumping arises due to appearance of intersite soliton solutions around $\theta=\pi$ (see red lines in b). See Supplementary Information Section S-5 for stability analysis at $\theta=\pi$. Here, we set $m_{0}=g=1$ and $N=1.45$.
Figure 2a illustrates that when $g_{12}=1$, the soliton’s center of mass travels across the lattice by 1-1 over one period, which is equal to the Chern number of the lowest band. This result is consistent with the previous findings [35, 36, 37, 38]. Surprisingly, we observe that when $g_{12}=0$, the soliton’s displacement changes to 0, 2-2, or 3-3, which differ from the Chern number of the lowest band (see also Figs. 2b-d for evolution of the soliton profiles). Here, we set $g=1$ (see Supplementary Information Section S-4 for discussion on effects of $g$ on soliton’s width).
[…]
@@
We proceed to examine how nonlinear Thouless pumping transitions to the anomalous one. To illustrate this, we plot in Fig. 3 the center-of-mass position of stable nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as a function of $\theta$ for different values of $g_{12}$. Notably, when $g_{12}=0.3$, the center-of-mass positions of a soliton and the Wannier centers exhibit clear differences for most values of $\theta$, but they coincide at $\theta=0$ and $\theta=\pi$. In fact, at $\theta=\pi$, the soliton solution is precisely the Wannier function multiplied by $\sqrt{N}$ when $m_{0}=1$ (see Supplementary Information Section S-5). However, as $g_{12}$ decreases, a stable intersite soliton solution emerges at $\theta=\pi$, localized around $j=l$ with $l$ being an integer, differing from the Wannier center, which is located at $l+1/2$. With a further decrease in $g_{12}$, we observe that the new branch of solutions becomes longer, eventually connecting to the original soliton solutions (represented by black lines). Consequently, when $g_{12}<0.2$, a soliton becomes displaced by 2-2 during adiabatic evolution.
@@
Nonlinearity-induced Thouless pumping.—We have demonstrated the breakdown of the correspondence between displacement and the Chern number for a nonlinear Thouless pump, which naturally leads to a question whether a soliton can be pumped when the Chern number of the linear band is zero. Our answer is affirmative. To illustrate this, we again consider the linear Hamiltonian in Eq. (2). Now, we vary the system parameters $m_{z}$, $J_{1}=J_{1}^{\prime}$ and $J_{2}$ with respect to $\theta$. For simplicity, we impose the conditions that all these parameters are symmetric about $\theta=\pi$ as illustrated in Fig. 4a. Consequently, the Hamiltonian $H^{\textrm{lin}}(k,\theta)$ and the Zak phase $\gamma_{n}(\theta)$ also have this symmetry; $\gamma_{n}(\theta)=i\int_{0}^{2\pi}{\rm d}k\langle u_{n,k}|\partial_{k}u_{n,k}\rangle$ is the Zak phase of the 1D system at a fixed $\theta$ with $|u_{n,k}\rangle$ being the $n$th eigenstate of $H^{\textrm{lin}}(k,\theta)$. The Chern number for each band is thus zero so that the linear Thouless pumping cannot occur.
@@
To achieve nonlinearity-induced Thouless pumping of solitons, we simultaneously adjust $g_{12}$, as shown in Fig. 4b (the interaction can be controlled via Feshbach resonances in ultracold atomic gases [25]). Specifically, we initially reduce $g_{12}$ to zero gradually, which facilitates the soliton’s travel over a longer distance. Subsequently, we slowly increase $g_{12}$ back to its original value to prevent the soliton from returning to its initial position. In Fig. 4c, we plot the density distribution of nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as $\theta$ varies from 0 to $2\pi$, remarkably demonstrating that the soliton is displaced by 1-1 unit cell. This effect is further evidenced in Fig. 4d, which depicts the evolution of center-of-mass positions of instantaneous soliton solutions as $\theta$ changes. This observation starkly contrasts with the behavior of the Wannier center (or the Zak phase), which returns to its starting value after one complete cycle. Hence, we successfully achieve Thouless pumping of a soliton through the influence of nonlinearity.
@@
Figure 4: Nonlinearity-induced Thouless pumping. a,b, Illustration of how the parameters $m_{z}(\theta)$, $J_{1}(\theta)$, $J_{2}(\theta)$ and $g_{12}(\theta)$ are varied with respect to $\theta$ in order to induce a nonlinear pumping for a topologically trivial band. c, One-cycle evolution of the density distribution $|\psi_{j}|^{2}$ of instantaneous solitons bifurcating from the lowest band. d, The evolution of center-of-mass positions of instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one cycle. Here, we set $N=1.25$ and $g=1$.